Tags:
create new tag
view all tags

20130124 Shift

AA, LN, KK

  • Recommission laser-wire
  • Signal correlations
  • Ideally collisions

  • Signal/Noise vs charge
  • e beam angle correlation
  • charge correlation
  • position correlation
  • reference MFB2FF optimal position
  • laser-wire absolute alignment reference

  • several scans, same e beam, different number of point
  • several scans, same e beam, different number of samples

Startup

  • Linac master oscillator is 1.428047000 GHz -> 357.01175. At least 10kHz different from last week. May need to tune seed.

Laser RGA not working since day before yesterday. Limited time before shift due to ATF2 project meeting. Build up time in RGA significantly longer (~200ns) than usual. Still 40ns difference between seeded and unseeded - definitely RGA on its own. Obvious, thing is flash lamp timing relative to pockel's cells then if not that, deal with physical alignment, starting with slight tweak to cavity mirror, then adjustment of etalon angle.

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r1 - 24 Jan 2013 - LaurieNevay

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback